A certain aircraft can fly 1190 miles with the wind in 5 hours and travel the same distance against the wind in 7 hours. What is the speed of the wind?
Answer by mananth(16949)

D= rt
wind rate = y
plane rate = x
distance =1190
time = 5 hours
1190/x+y=5
5x+5y=1190
/5
x+y=238.................1
...
against wind
1190/(x-y)=7
7x-7y=1190
x-y = 170................2
add the two equations
2x=408
x= 204 mph plane speed
238-204=34 mph speed of wind